Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Mon Nov 30 10:36:52 MET 1998) Success
No problems except a few seconds earlier stops of tape by the low tape sensors.

Medicina (Venturi T. at Mon Nov 30 11:26:34 MET 1998) Minor failures
No apparent problems recorded during the experiment.

Medicina (Venturi T. at Mon Nov 30 11:27:01 MET 1998) Success
No apparent problems recorded during the experiment.

Effelsberg (dgraham at Mon Nov 30 15:18:56 MET 1998) Minor failures
For unknown reasons, data flagged as bad 0700-1000 and 1200-1215. Tcal values also different in LCP and RCP.

Jodrell (Alastair Gunn at Tue Dec 1 08:14:12 MET 1998) Minor failures
Missed scans from 8:08 to 9:20 due to PI not giving adequate slew time for the Lovell telescope. Short lived Servo failure at approx 13:07. No other known problems.

Westerbork (Tony Foley at Tue Dec 1 18:36:47 MET 1998) Minor failures
Vacuum lost on tapes at 11:40 and 13:58UT, with heavy deposits of dirt coming off the tape; however for much of this time (12:57 - 14:10) the scans were out of Hour Angle range anyway. There will probably be problems reading this tape.

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:45:10 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gp019.antab. Read the comments in the file to get further information.